# Large Language Models Synergize with Automated Machine Learning

### v1

@@ 3.1 Decomposition and compatibility
Compared to typical sequence generation tasks, code generation demands a complex output sequence. For example, generating 10 lines of code to meet a certain requirement is often considerably more challenging than an open-ended short writing of similar length. In addition, for an ML pipeline, certain components—such as data preparation, modeling, and post-processing—require task-specific solutions. The components constitute a lengthy concatenated sequence, which can be multiple times longer than those found in traditional coding problems.
@@
This framework is a form of compositional reasoning Andreas et al. (2016), which solves challenging problems by breaking them down into smaller, manageable components or modules. The framework also mirrors the intuitive practices of many human programmers who break down a large program into manageable chunks, verifying the functionality and compatibility of each segment during its creation. While the framework is generic for any sequence generation tasks under the availability of sequence evaluation methods defined in Definition 3.1, the framework is particularly effective in generating ML programs.
Figure 5: Contexual Modular Generation for ML tasks. First, the task description is given to the LLM. Then the LLM generates each module (i.e., the code for a step in ML) sequentially. Each module is generated in an iterative process, which includes the generation of code by the LLM, followed by verification via the evaluation methods (unit tests in Section 4), and incorporating feedback from these evaluations back into the LLM. In the end, the generated modules are automatically integrated into the code of the ML pipeline.
@@ 3.2 Formal definition
We consider a scenario where a binary evaluation method, denoted as $\mathbb{E}$, is available to test whether a generated sequence meets a specified requirement (e.g., absence of detectable errors and presence of desired functionality). For a given task, there exists a set of sequences that fulfill this requirement, represented as $\mathbf{\Omega}=(\omega_{1},\omega_{2},\ldots,\omega_{C})$. Let $\mathbb{E}(Y)$ denote that the sequence $Y$ has been assessed by $\mathbb{E}$ and meets the specification. To establish a fair comparison framework applicable to all sequence generation methods (refer to Sections 2.1 and A.1), we focus on the task of repeatedly generating the target sequence until a satisfactory sequence is produced. This task involves an LLM with state $\lambda$ and a set of input sequences ${X_{i}}$. In each $i_{th}$ iteration, the LLM generates an output sequence $Y_{i}$ using $X_{i}$. The process continues until the LLM produces a sequence $Y$, such that $Y$ can be any sequence or any concatenation of sequences generated in the current or previous iterations that meet the user’s requirements as verified by $\mathbb{E}$.
@@
1:Input: An autoregressive LLM $f_{\lambda}$, number of modules $J$, initial instructions for generating each module $\{X^{1}_{0},X^{2}_{0},\ldots,X^{J}_{0}\}$, contextual modular evaluations for each module $\{\mathbb{E}^{1},\mathbb{E}^{2},\ldots,\mathbb{E}^{J}\}$, and instruction constructors for each module $\{\Phi^{1},\Phi^{2},\ldots,\Phi^{J}\}$
@@
Algorithm 1 defines our sequence generation framework (Figure 5). As a form of compositional reasoning, Algorithm 1 first divides the entire program into $J$ modules. Then for each module $j$, the algorithm repeatedly generates the code until success (lines 4 to 11). In the $i_{th}$ iteration, the LLM $f_{\lambda}$ first generates the code $Y^{j}$ using the instruction $X^{j}_{i}$ (line 8). Then, the evaluation method for the module $j$, $\mathbb{E}^{j}$, examines $Y^{j}_{i}$ and produce the feedback $R_{i}$ (line 9) . In line 6, previous instructions $X^{j}_{<i}$, generated code $Y^{j}_{<i}$, and feedbacks $R_{<i}$ are assembled as the instruction $X^{j}_{i}$ for the current iteration.
@@
For a task with desired sequences $\mathbf{\Omega}$, suppose the output sequences in one decomposition are $\{Y^{1},Y^{2},\ldots,Y^{J}\}$. For a set of binary evaluation methods $\{\mathbb{E}^{1},\mathbb{E}^{2},\ldots,\mathbb{E}^{J}\}$. We call the set of binary evaluation methods as contextual modular evaluations for the task, if that every $Y^{j}$ in $\{Y^{1},Y^{2},\ldots,Y^{J}\}$ is verified as valid by the corresponding $\mathbb{E}^{j}$ in $\{\mathbb{E}^{1},\mathbb{E}^{2},\ldots,\mathbb{E}^{J}\}$ garrantees that $\{Y^{1},Y^{2},\ldots,Y^{J}\}$ form a valid prgram (if $((\forall j,\mathbb{E}^{j}(Y^{j}))\rightarrow\{Y^{1},Y^{2},\ldots,Y^{J}\}\in\mathbf{\Omega}))$).
In other words, if the generated modules are verified by the contextual modular evaluations, then the modules are compatible with each other to form a desired program without any modification. Contextual modular evaluations eliminate the need for extra computation for adapting the generated modules into a sequence $Y$ such that $Y\in\mathbf{\Omega}$. In our study, the contextual modular evaluations verify individual components in the ML pipeline to ensure consistency in data flow based on a novel testing technique (Section 4).
@@ 3.3 Theoretical analysis
Next, we theoretically analyze the scalability of Algorithm 1 over the length of complex output sequences. First, we define several metrics for comparing sequence generation methods. Under the comparison scheme defined in Section 3.2, suppose the process of repeated generation terminates and the total number of iterations is $G$ (We count $G$ as $\infty$ if the iteration never terminates). Then the average number of generations per success $\mathbf{E}(G)$ is a measurement of the generation efficiency. There are multiple ways to define the output length of a sequence generation problem Anil et al. (2022). In our study, we define the output length of the task as the length of the longest valid output sequence: $\Gamma=max\{|Y||Y\in\mathbf{\Omega}\}$.
For a sequence $\alpha$ and an integer $\beta$, let $\Theta_{\Omega}(\alpha,\beta)$ denote the set of sequences $\{s|\exists s_{r}\land\exists\gamma(1\leq\gamma\leq\Gamma),|s|=\beta\land|(\alpha,s,s_{r})|=\gamma\land(\alpha,s,s_{r})\in\mathbf{\Omega}\}$, where $s_{r}$ is an arbitrary sequence (including the empty sequence) and $|\alpha|+\beta\leq\Gamma$. Intuitively, in a sequence generation process, $\Theta_{\Omega}(\alpha,\beta)$ represents all valid sub-sequences from the ${(|\alpha|+1)}_{th}$ position to the ${(|\alpha|+\beta)}_{th}$ position in the output sequence, given that the previously generated part in the output sequence is $\alpha$. If $\beta=1$, then $\Theta_{\Omega}(\alpha,\beta)$ represents all valid tokens at the ${(|\alpha|+1)}_{th}$ position. For example, in Figure 6, let $t_{i}$ denotes the $i_{th}$ candidate token, then $\Theta_{\Omega}((t_{2},t_{4},t_{6}),1)=\{t_{1},t_{3}\}$.
@@
For any sequence generation process, if there exists a step in the process that generates the entire output sequence and Assumption 3.4 holds for the step, then the best case complexity of $\mathbf{E}(G)$ of the process is $\mathcal{O}(exp(\Gamma))$.
@@
Assumption 3.4 posits that errors may occur for any token of the output sequence, and the validity of this assumption is contingent upon $\Omega$, $\lambda$, and $X$. Theorem 3.5 suggests that for traditional sequence generation approaches based on autoregressive LLMs, if the task exceeds a certain complexity level, the generation difficulty escalates exponentially over the length of the problem. If the value of $\xi_{X}^{\Omega,\lambda}(\epsilon)$ is ascertainable or estimable for a given $\epsilon$, regardless of the applicability of Assumption 3.4, deriving a more specific lower bound of $\mathbf{E}(G)$ might be possible (Appendix B).
@@
If Assumption 3.7 holds, then for Contextual Modular Generation as defined in Algorithm 1, the best case complexity of $\mathbf{E}(G)$ is $\mathcal{O}(\Gamma)$.
@@ 4 Unit Tests for Machine Learning
Contextual Modular Generation is conceptually intuitive, but its implementation demands meticulous design and non-trivial efforts. In this study, we utilize unit tests as contextual modular evaluations (as defined in Definition 3.1). The desiderata for the unit tests in our study are:
@@
  • Internal compatibility: In our method, for each module in the ML pipeline, multiple code implementations of different candidate methods are generated to construct multiple candidate pipelines. (Figure 7 and Figure 3). Different verified candidate codes of any module should be compatible with the candidate codes of other modules. This compatibility is crucial for the unit tests’ role as contextual modular evaluations (as defined in Definition 3.1). Furthermore, for a pipeline with $N$ modules, if $M$ candidate codes for each module are generated, this compatibility increases the ratio of generated pipeline combinations over generation cost by a factor of up to $M^{N-1}$ (Figure 7).
  • External compatibility: Verified modules should also be compatible with pre-written modules, including those implementing a consistent evaluation procedure, training procedure, validation procedure, and autoML methods (Figure 7).
@@
Figure 7: Constrained Generative Unit Testing. The yellow arrow represents the data flow between codes. The unit tests focus on verifying the structure and content of the data exchanged among the codes. For each module, codes for different candidate methods are generated. With the pre-written modules, each pathway through the network forms a pipeline.
@@ 5 Optimization of Machine Learning Programs
1:Input: An autoregressive LLM $f_{\lambda}$ and textual description of an ML task $\mathbb{T}$
$\mathbb{H}=[\;]$ and $\mathbf{S}=Search\;Space\;Generation\;(\mathbb{T},f_{\lambda})$
3:for $\theta$ in $\mathbf{S}$ do
@@
6:if based on $\mathbf{S}$, the task is a deep learning task then
7:     $\mathbf{S}=ZC\;Proxies\;Evaluation\;(\mathbf{S},\{\mathbf{C}^{m}_{\theta}\})$
@@
10:     $\theta=Search\;Strategy\;(\mathbf{S},\mathbb{H})$
11:     for non-modeling module $\mathbb{M}$ in $\theta$ do
12:         if no prior attempt exists for generating $\mathbf{C}^{\theta}_{\mathbb{M}}$ then
13:              $\mathbf{C}^{\theta}_{\mathbb{M}}=Module\;Generation\;(\mathbb{M},\theta,f_{\lambda})$
@@
17:         Evaluate the ML program based on $\theta$, with evaluation result as $r_{\theta}$, then append $[\theta,r_{\theta}]$ to $\mathbb{H}$
@@
As Figure 7 illustrates, a network of codes is generated to form multiple candidate pipelines. AutoML then selects the high-performance pipeline (i.e., pathway within the network), along with associated hyperparameter values. With the external compatibility (Section 4), autoML is implemented as pre-written modules.
Many autoML methods architect high-performance deep learning models from scratch—such as determining the number of convolutional layers and initiating training from zero. Such a process is prohibitively expensive for regular users under low resource constraints Öztürk et al. (2022). By comparison, the search space in our study leverages the ability of LLMs to directly deploy pretrained models as candidate methods for modeling.
@@
Algorithm 2 outlines our comprehensive methodology, from code generation to performance optimization, as illustrated in Figure 3. In Algorithm 2, the LLM $f_{\lambda}$ initially creates the search space $\mathbf{S}$ with Search Space Generation (Appendix C.2), suggesting candidate methods for each module, such as candidate feature engineering techniques, ML models, and hyperparameter values (line 2). Then for the ML model in each candidate solution of the search space $\theta$, generates the corresponding code for the model $\mathbf{C}^{\theta}_{\mathbf{M}}$ with Module Generation (Appendix C.4) (lines 4 to 6). Module Generation generates the code with the LLM based on Algorithm 1 and Algorithm 3.
Notably, with the synthetic data from Algorithm 3, for deep learning tasks, the modeling modules can be evaluated with ZC Proxies Evaluation (Appendix C.6) before generating other modules, saving both generation and optimization costs (lines 7 to 9). ZC Proxies Evaluation removes a fraction of lower-performing candidate solutions from the search space.
Subsequently, Algorithm 2 iteratively examines the candidate solutions in the search space through trial-and-error. In each iteration, with the previous optimization history $\mathbb{H}$ and search space $\mathbf{S}$, Search Strategy (Appendix C.3), which is an optimization algorithm, selects a promising solution for the ML task (line 10). Then, for the selected solution, Module Generation generates all modules except the modeling module (lines 11 to 15). Once all valid modules for the solution are created, Algorithm 2 assembles the complete program for the ML pipeline, evaluates it using the designated numerical metrics in $\mathbb{T}$, and records the solution and its performance in the optimization history.

### v2

@@ 3.1 Decomposition and compatibility
Compared to typical sequence generation tasks, code generation demands a complex output sequence. For example, generating 10 lines of code to meet a certain requirement is often considerably more challenging than an open-ended short writing of similar length. In addition, for an ML program, certain components—such as data preparation, modeling, and post-processing—require task-specific solutions. The components constitute a lengthy concatenated sequence, which can be multiple times longer than those found in traditional coding problems.
@@
This framework is a form of compositional reasoning (Andreas et al., 2016), which solves challenging problems by breaking them down into smaller, manageable components or modules. The framework also mirrors the intuitive practices of many human programmers who break down a large program into manageable chunks, verifying the functionality and compatibility of each segment during its creation. While the framework is generic for any sequence generation tasks under the availability of sequence evaluation methods defined in Definition 3.1, the framework is particularly effective in generating ML programs.
Figure 6: Contexual Modular Generation for ML tasks. First, the task description is given to the LLM. Then the LLM generates each module (i.e., the code for a step in ML) sequentially. Each module is generated in an iterative process, which includes the generation of code by the LLM, followed by verification via the evaluation methods (unit tests in Section 4), and incorporating feedback from these evaluations back into the LLM. In the end, the generated modules are automatically integrated into the ML program.
@@ 3.2 Formal definition
Consider the scenario where a binary evaluation method, denoted as $\mathbf{E}$, is available to test whether a generated sequence meets a specified requirement (e.g., absence of detectable errors and presence of desired functionality). For a given task, there exists a set of sequences that fulfill this requirement, represented as $\mathbf{\Omega}=(\omega_{1},\omega_{2},\ldots,\omega_{C})$. Let $\mathbf{E}(Y)$ denote that the sequence $Y$ has been assessed by $\mathbf{E}$ and meets the specification. To establish a fair comparison framework applicable to all sequence generation methods (see Section 2.1 and Appendix A.1), we focus on the task of repeatedly generating the target sequence until a satisfactory sequence is produced. This task involves an LLM with state $\lambda$ and a set of input sequences ${X_{i}}$. In each $i_{th}$ iteration, the LLM generates an output sequence $Y_{i}$ using $X_{i}$. The process continues until the LLM produces a sequence $Y$, such that $Y$ can be any sequence or any concatenation of sequences generated in the current or previous iterations that meet the user’s requirements as verified by $\mathbf{E}$.
@@
1:Input: An autoregressive LLM $f_{\lambda}$, number of modules $J$, initial instructions for generating each module $\{X^{1}_{0},X^{2}_{0},\ldots,X^{J}_{0}\}$, contextual modular evaluations for each module $\{\mathbf{E}^{1},\mathbf{E}^{2},\ldots,\mathbf{E}^{J}\}$, and instruction constructors for each module $\{\Phi^{1},\Phi^{2},\ldots,\Phi^{J}\}$
@@
Algorithm 1 defines our sequence generation framework (Figure 6). As a form of compositional reasoning, Algorithm 1 first divides the entire program into $J$ modules. Then for each module $j$, the algorithm repeatedly generates the code until success (lines 4 to 11). In the $i_{th}$ iteration, the LLM $f_{\lambda}$ first generates the code $Y^{j}$ using the instruction $X^{j}_{i}$ (line 8). Then, the evaluation method for the module $j$, $\mathbf{E}^{j}$, examines $Y^{j}_{i}$ and produce the feedback $R_{i}$ (line 9) . In line 6, previous instructions $X^{j}_{<i}$, generated code $Y^{j}_{<i}$, and feedbacks $R_{<i}$ are assembled as the instruction $X^{j}_{i}$ for the current iteration.
@@
For a task with desired sequences $\mathbf{\Omega}$, suppose the output sequences in one decomposition are $\{Y^{1},Y^{2},\ldots,Y^{J}\}$. For a set of binary evaluation methods $\{\mathbf{E}^{1},\mathbf{E}^{2},\ldots,\mathbf{E}^{J}\}$. We call the set of binary evaluation methods as contextual modular evaluations for the task, if that every $Y^{j}$ in $\{Y^{1},Y^{2},\ldots,Y^{J}\}$ is verified as valid by the corresponding $\mathbf{E}^{j}$ in $\{\mathbf{E}^{1},\mathbf{E}^{2},\ldots,\mathbf{E}^{J}\}$ garrantees that $\{Y^{1},Y^{2},\ldots,Y^{J}\}$ form a valid prgram (if $((\forall j,\mathbf{E}^{j}(Y^{j}))\rightarrow\{Y^{1},Y^{2},\ldots,Y^{J}\}\in\mathbf{\Omega}))$).
In other words, if the generated modules are verified by the contextual modular evaluations, then the modules are compatible with each other to form a desired program without any modification. Contextual modular evaluations eliminate the need for extra computation for adapting the generated modules into a sequence $Y$ such that $Y\in\mathbf{\Omega}$. In our study, the contextual modular evaluations verify individual components in the ML program to ensure consistency in data flow based on a novel testing technique (Section 4).
@@ 3.3 Theoretical analysis
Next, we theoretically analyze the scalability of Algorithm 1 over the length of complex output sequences. First, we define several metrics for comparing sequence generation methods. There are multiple ways to define the output length of a sequence generation problem (Anil et al., 2022). In our study, we define the output length of the task as the length of the longest valid output sequence: $\Gamma=max\{|Y||Y\in\mathbf{\Omega}\}$. Under the comparison scheme defined in Section 3.2, suppose the process of repeated generation terminates and the total number of iterations is $G(\Gamma)$ (We count $G(\Gamma)$ as $\infty$ if the iteration never terminates). Then the expectation of generations per success $\mathbb{E}_{x\sim{P}}[G(\Gamma)]$ is a measurement of the generation efficiency.
For a sequence $a$ and an integer $n$, let $\Theta_{\Omega}(a,n)$ denote the set of sequences $\{s|\exists s_{r}\land(\exists m\;(1\leq m\leq\Gamma),|s|=\beta)\land|(a,s,s_{r})|=m\land(a,s,s_{r})\in\mathbf{\Omega}\}$, where $s_{r}$ is an arbitrary sequence (including the empty sequence) and $|a|+n\leq\Gamma$. Intuitively, in a sequence generation process, $\Theta_{\Omega}(a,n)$ represents all valid sub-sequences from the ${(|a|+1)}_{th}$ position to the ${(|a|+n)}_{th}$ position in the output sequence, given that the previously generated part in the output sequence is $a$. If $n=1$, then $\Theta_{\Omega}(a,n)$ represents all valid tokens at the ${(|a|+1)}_{th}$ position. For example, in Figure 7, let $t_{i}$ denotes the $i_{th}$ candidate token, then $\Theta_{\Omega}((t_{2},t_{4},t_{6}),1)=\{t_{1},t_{3}\}$.
@@
For any sequence generation process, if there exists a step in the process that generates the entire output sequence and Assumption 3.4 holds for the step, then the best case complexity of $\mathbb{E}_{x\sim{P}}[G(\Gamma)]$ of the process is $\mathcal{O}(exp(\Gamma))$.
@@
Assumption 3.4 posits that errors may occur for any token of the output sequence, and the validity of this assumption is contingent upon $\mathbf{\Omega}$, $\lambda$, and $X$. Theorem 3.5 suggests that for traditional sequence generation approaches based on autoregressive LLMs, if the task exceeds a certain complexity level, the generation difficulty escalates exponentially over the length of the problem. If the value of $\xi_{X}^{\Omega,\lambda}(\epsilon)$ is ascertainable or estimable for a given $\epsilon$, regardless of the applicability of Assumption 3.4, deriving a more specific lower bound of $\mathbb{E}_{x\sim{P}}[G(\Gamma)]$ might be possible (Appendix B).
@@
If Assumption 3.7 holds, then for Contextual Modular Generation as defined in Algorithm 1, the best case complexity of $\mathbb{E}_{x\sim{P}}[G(\Gamma)]$ is $\mathcal{O}(\Gamma)$.
@@ 4 Unit Tests for Machine Learning
Contextual Modular Generation is conceptually intuitive, but its implementation demands meticulous design and non-trivial efforts. In this study, we utilize unit tests as contextual modular evaluations (as defined in Definition 3.1) to ensure compatibility among the modules. The desiderata for the unit tests in our study are:
@@
  • Internal compatibility: In our method, for each module in the ML program, multiple code implementations of different candidate methods are generated to construct multiple candidate programs. (Figure 8 and Figure 3). Different verified candidate codes of any module should be compatible with the candidate codes of other modules. This compatibility is crucial for the unit tests’ role as contextual modular evaluations (as defined in Definition 3.1). Furthermore, for an ML program with $N$ modules, if $M$ candidate codes for each module are generated, this compatibility increases the ratio of generated program combinations over generation cost by a factor of up to $M^{N-1}$ (Figure 8).
  • External compatibility: Verified modules should also be compatible with pre-written modules, including those implementing a consistent evaluation procedure, training procedure, testing procedure, and autoML methods (Figure 8).
@@
Figure 8: Constrained Generative Unit Testing for CV tasks. The yellow arrow represents the data flow between codes. The unit tests focus on verifying the structure and content of the data exchanged among the codes. For each module, codes for different candidate methods are generated. With the pre-written modules, each pathway through the network forms an ML program.
@@ 5 Optimization of Machine Learning Programs
1:Input: An autoregressive LLM $f_{\lambda}$ and textual description of an ML task $\mathbf{T}$
$\mathbb{H}=[\;]$ and $\mathbb{S}=Search\;Space\;Generation\;(\mathbf{T},f_{\lambda})$
3:for $\theta$ in $\mathbb{S}$ do
@@
6:if based on $\mathbb{S}$, the task is a deep learning task then
7:     $\mathbf{S}=ZC\;Proxies\;Evaluation\;(\mathbb{S},\{\mathbf{C}^{m}_{\theta}\})$
@@
10:     $\theta=Search\;Strategy\;(\mathbb{S},\mathbb{H})$
11:     for module $j$ in $\theta(j\neq\mathbf{M})$ do
12:         if no prior attempt exists for generating $\mathbf{C}^{\theta}_{j}$ then
13:              $\mathbf{C}^{\theta}_{j}=Module\;Generation\;(j,\theta,f_{\lambda})$
@@
17:         Evaluate the ML program based on $\theta$, with evaluation result as $R_{\theta}$, then append $\theta$ and $R_{\theta}$ to $\mathbb{H}$
@@
As Figure 8 illustrates, a network of codes is generated to form multiple candidate ML programs. AutoML then selects the high-performance program (i.e., pathway within the network), along with associated hyperparameter values. With the external compatibility (Section 4), autoML is implemented as pre-written modules.
Many autoML methods architect high-performance deep learning models from scratch—such as determining the number of convolutional layers and initiating training from zero. Such a process is prohibitively expensive for regular users under low resource constraints (Öztürk et al., 2022). By comparison, the search space in our study leverages the ability of LLMs to directly deploy pretrained models as candidate methods for modeling.
@@
Algorithm 2 outlines our comprehensive methodology, from code generation to code optimization, as illustrated in Figure 3. In Algorithm 2, the LLM $f_{\lambda}$ initially creates the search space $\mathbb{S}$ with Search Space Generation (Appendix C.2), suggesting candidate algorithms for each module, such as candidate feature engineering techniques and ML models, and candidate hyperparameter values (line 2). Then for the ML model in each candidate solution of the search space $\theta$, generates the corresponding code for the model $\mathbf{C}^{\theta}_{\mathbf{M}}$ with Module Generation (Appendix C.4) (lines 4 to 6). Module Generation generates the code with the LLM based on Algorithm 1 and Algorithm 3.
Notably, with the synthetic data from Algorithm 3, for deep learning tasks, the modeling modules can be evaluated with ZC Proxies Evaluation (Appendix C.6) before generating other modules, saving both generation and optimization costs (lines 7 to 9). ZC Proxies Evaluation removes a fraction of lower-performing candidate solutions from the search space with negligible costs.
Subsequently, Algorithm 2 iteratively examines the candidate solutions in the search space through trial-and-error. In each iteration, with the previous optimization history $\mathbb{H}$ and search space $\mathbb{S}$, Search Strategy (Appendix C.3), which is an optimization algorithm, selects a promising solution for the ML task (line 10). Then, for the selected solution, Module Generation generates all modules except the modeling module (lines 11 to 15). Once all valid modules for the solution are created, Algorithm 2 assembles the complete ML program, evaluates it using the designated numerical metrics in $\mathbf{T}$, and records the solution and its performance in the optimization history.